• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 16:00:24 on May19,2007

    Entry number 202692

    keyword=day shift summary

    08:00-: smooth running. Occasional Left Q3 drifts fixed by changing P0 Set value by 1E-5 GeV/c. Good beam

    10:30 Error message with 3630 end-of-run, see halog entry 202664. Also the start-of-run entry showed up after run ends, probablly related to the same error. No events summary in end-of-run entry but data look okay (checked with replays).

    10:30 Move target to Al dummy. Took a very short run #3631 to check prescaling and make sure we don't have the same error as 3630 (and it didn't).

    10:40-11:40 Taking data with Al 20cm dummy, 40uA beam, run 3632.

    11:42 Back to 3He target. Smooth running...
    15:00 ABU this shift: 7.04 hours!!!

    Good 3He runs:
    3628, 1.25M
    3629, 1.25M
    3630, 1.25M, error message at end-of-run summary.
    3633, 1.25M
    3634, 1.25M
    3635, 1.25M
    3636, 1.25M
    3637, 1.25M

    Good Al dummy runs:
    3632, 750K